# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v12

@@ B.6.2 Ancillary Wage
Notes: This scatter plot shows the relationship between the log-transformed number of factory workers and the log-transformed share of ancillary wage. The raw data are listed in Table B.3. The dashed line indicates the linear fit using the log-transformed variables: the fitted regression is $\hat{y}=0.747+0.428x$ with $R^{2}=0.82$. The predictions are materially similar if I use the fractional polynomial using the raw data on the number of factory workers and share of ancillary wage in Table B.3. In the case of fitting based on the raw data, the fitted fractional polynomial model is $\hat{y}^{\textit{Daily Wage}}=53.6-399.3a^{-1/2}+(2.84\times 10^{-10})a^{3}$ with $R^{2}=0.87$. I prefer to use linear regression herein because it can provide more precise fitted values for small-and medium-scale factories.
Source: Kitazawa (1924, pp. 20; 58–62).
The daily wage estimated in Section B.6.1 is a fixed salary and thus, does not include additional salaries such as bonuses and allowances (hereafter, wages other than regular salaries are collectively referred to as ancillary wages). However, there are no systematic statistics on the ancillary wages throughout the prewar period (Committee for the History of the Labor Movement 1964, pp. 14–15). Moreover, since the Tsukishima Survey targets skilled worker households, information on the ancillary wages paid to skilled workers is necessary. Fortunately, I found a survey report named “Wage Survey Report” edited by Kitazawa (1924). This report includes the result of a survey on the wages for machine finishers (fitters) and turners at six machine factories in Tokyo from May–October 1921. Thus, this is particularly useful material to know the wage structure of skilled factory workers.
@@
Panel A in Table B.3 summarizes the survey results. The surveyed factories included $86$, $157$, $376$, $623$, $928$, and $2,837$ workers, ranging from small to large plants. The percentage of skilled workers (finishers and turners) in the surveyed factories ranged from $27$% to $77$%, and the portion of monthly ancillary wages was 12–53%. Figure B.9 is a linear regression fit using log-transformed data on the percentage of ancillary wages and the number of workers. There is a positive correlation between the factory size and ancillary wages. This is consistent with the historical fact that larger factories have more extensive benefits than smaller ones (Odaka 1999).
@@
Panel B of Table B.3 summarizes the distributions of the number of factories, the total number of workers, and the estimated average share of ancillary wage in Tokyo City (Column 1) and Tsukishima (Column 2). The average ancillary wage at the Tsukishima factories is $12$%, which is slightly larger than the value for the entire city of Tokyo ($9$%). The percentage shares of the number of factory workers by each bin are listed in Columns (1) and (2) of Panel C in Table B.3. As shown, Tsukishima factories tended to have larger factories than those of the entire city, which is reflected in the slightly larger estimate in the average ancillary wage for the Tsukishima factories. In Column (3) of Panel C in Table B.3, the number of THBS heads whose occupation, factory name, and factory size are available are listed. While the percentage shares of the heads in each bin show a similar distribution to that of the Tsukishima factories, the share of middle (large) sized factories is larger (smaller) than the population. If interpreted literally, this tendency implies that the (unobservable) true average ancillary wage for the THBS heads may be slightly lower than the population mean. However, note that the purpose of Panel C in Table 2 is to compare the average monthly wage of the skilled factory workers employed in the Tsukishima factories to the THBS heads’ average monthly earnings. Therefore, the purpose herein is to estimate the average ancillary wage for the target population, for example, the Tsukishima factories, but not to calculate the average ancillary wage for the THBS sample itself.
@@
Notes:
Panel A: This panel shows the size of factories and the share of fixed and ancillary monthly wages in six factories surveyed in Kitazawa (1924). The number of factory workers surveyed in factories A–F are 14, 23, 9, 11, 11, and 16, respectively. The share of these surveyed skilled workers to all skilled workers is 20.3 in Factory A, 50.0 in Factory B, 5.3 in Factory C, 2.3 in Factory D, and 4.2 in Factory F, respectively. This figure for factory E is unavailable because the number of skilled workers is unknown (Kitazawa 1924, p. 21).
Panel B: Subcolumns 1–2 and 4–5 summarize the number of factories and total number of workers in Tokyo City and Tsukishima, respectively (Handbook of Factories 1921). Subcolumns 3 and 6 show the average share of ancillary monthly wage predicted in Figure B.9 for Tokyo City and Tsukishima, respectively.
Panel C: Columns 1–3 summarize the number of factory workers and its share in percentage points in Tokyo City, Tsukishima, THBS sample, respectively. The statistics in Columns 1 and 2 (same figures used in Panel B) are from the Handbook of Factories (1921). In Column 3, the THBS heads who worked in the machinery factories are considered. There are 28 heads whose factory names and size are available. For a blacksmith included in these heads, I assume the number of workers to be ten, based on the report (Department of Health, Ministry of the Interior 1923a, p. 412).
Sources: Panel A: Kitazawa (1924, pp. 20; 58–62). Panel B: The number of factory workers is taken from the Ministry of Agriculture and Commerce (1921, pp. 613–629; 674). The average shares of ancillary wage are calculated using Kitazawa (1924, pp. 20; 58–62). See also Figure B.9. Panel C: Data on the number of factory workers and factories are from the Ministry of Agriculture and Commerce (1921, pp. 613–629; 674). Data on the number of heads are from the THBS sample.
@@ B.6.3 Estimating Monthly Income of Adult Male Factory Workers
Panel B: Estimated monthly earning: $w^{\textit{Total}}$
@@
Notes:
1. Average daily wage ($w^{Daily}_{j}$): The average daily wage of male factory workers in the skilled age range (aged 30–40) in the machinery sector is calculated using two official reports of the manufacturing censuses documented in the ASCT and LST. Details in the calculation steps are summarized in Online Appendix B.6.1.
@@
where $w^{\textit{Regular}}>0$ is the monthly regular wage and $w^{\textit{Ancillary}}\geq 0$ indicates the monthly ancillary wage. Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$.101101101Strictly, this weight may not be required to be in this range. I use this range because all the cases reported in Kitazawa (1924) take the weights between zero and one. The monthly wage equation is then specified as follows:
@@ B.7 Measurement Errors in the Unadjusted Panel Dataset
Note: Figures B.10aB.10cB.10e show the average semi-monthly income and expenditure calculated from unadjusted, shifted, and adjusted panel datasets, respectively. Unadjusted series defines the days from the 1st to the 15th of each month as the first half, whereas the days from the 16th to the end of each month as the second half. The shifted series defines the days from the 2nd to the 16th of each month as the first half, whereas the days from the 17th to the 1st of the following month are the second half. Adjusted series is the baseline definition in the main text, which uses the days from 12th to 26th in each month as the “first-half” and those from 27th to 11th in the next month as the “second-half.”
As explained in Section 3, the raw series contains considerable measurement errors due to the timing of payday. The shifted series can manage the measurement errors from the timing of payday but cannot adjust the difference in the payday timing and expenses. In other words, the shifted series ignores the households’ decision on the consumption schedule in each semi-month. The adjusted series reflecting the exact timing of the household expenses can manage measurement errors. Figures B.10bB.10dB.10f illustrate the average semi-monthly income minus expenditure calculated from unadjusted, shifted, and adjusted panel datasets, respectively.
Source: Created by the author using the THBS sample.
@@ B.8 Testing Attenuation Bias in the Unadjusted Panel Dataset
In this subsection, I assess the validity of the aggregation method used in this study. Adjusting the aggregate data to account for the timing of paydays reveals more transparent relationships between income and expenditure than the unadjusted data (Section 3.3). To demonstrate this mechanism, I listed the estimates from the unadjusted monthly panel dataset in Column (2) of Table B.5. The estimate for total consumption is $0.26$, which is much lower than the estimate for the adjusted series ($0.43$) in Column (1). In addition, most subcategories show systematically smaller estimates than those from the adjusted data listed in Column (2). For example, the estimate for the food category is now $0.10$ and statistically insignificant. This means that measurement errors in the unadjusted aggregation lead to strong attenuation bias, which results in false negatives (i.e., false rejections of the null of full risk-sharing hypothesis) in most of the consumption categories. I, therefore, specified a generating process of systematic measurement errors in the aggregate survey data, which has attracted attention in the macroeconomics literature (Cochrane 1991; Dynarski et al. 1997; Gervais and Klein 2010; Nelson 1994). Moreover, several subcategories select values close to one or large negative values, which are no longer economically interpretable.102102102For example, the estimate for housing was negative. The households paid a fixed rent amount monthly. Given that the rent payday was fixed, the shifts in wage paydays might create cells with less income with a fixed amount of rent. In the case of seasonal consumption categories, the shifts may coincide with the timing of seasonal spending. The THBS households tended to purchase work clothes in March to prepare for the start of the new fiscal year in April. They also spent more on clothing in June, likely to buy summer clothes. The shift of paydays to these months may have created a strong but false positive correlation. This means that the miss-assignment of paydays in the unadjusted data may cause critical statistical inference issues in a few subcategories.
@@ B.9 Considering Child Share as a Preference Shock Variable
As explained in the main text, the child share variable exhibits significantly more within-variations than that of the family size. Table B.6 presents the results from specifications that include the share of children under the primary school age range, specifically those aged 0–5, as an alternative preference shock variable. Panels A and B show the results for the consumption smoothing and risk-coping strategy regressions, respectively. The results are materially similar to the main results presented in the main text (Tables 4 and 5). Note that the results remain robust to different definitions of age bins (not reported).
@@
(1) Semi-monthly Panels (2) Adjusted Monthly Panels
@@
(1) Semi-monthly Panels (2) Adj. Monthly Panels
@@
Other member’s earnings 0.001 [0.010] 289 0.036 [0.033] 124
@@
Notes: Panel A: This table shows the results of Equation 1. The regressions of the 11 measures of log-transformed consumption on log-transformed disposable income, the proportion of children aged 0–5, household-fixed effect, and time-fixed effect are listed. Columns 1 and 2 show the results for the semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on log-transformed disposable income are listed in the columns named “Coef.”.
Panel B: This table shows the results of equation 2. The proportion of children aged 0–5, household fixed effects, and time-fixed effects are included in all the regressions. Columns 1 and 2 show the estimates for the semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on the head’s earnings are listed in the “Coef.” columns.
Panel B-1: Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts.
Panel B-2: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of miscellaneous assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
@@ B.10 Distribution of the Idiosyncratic Shocks
Notes: This figure shows the distribution of idiosyncratic shocks in the semi-monthly THBS panel dataset. The idiosyncratic shocks are estimated as the residuals from the regression of the first difference in income on the first difference in aggregate income. The red line indicates the mean of the idiosyncratic shocks ($-0.235$).
Source: Created by the author using the THBS sample.
Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset.103103103I use the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets. This is because the daily panels include a set of consecutive household-date cells with no income. These cells cannot be used to predict the idiosyncratic shocks as the residuals are economically meaningfully, and thus it is practically impossible to illustrate the idiosyncratic shocks based on the daily dataset. The idiosyncratic shocks are calculated as residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919. For household $i$ in semi-month $t$ cell, the residual is then defined as follows:
@@ B.11 Unit Root and Autocorrelations
Table B.7 presents the results of the Fisher-type panel unit root tests for the total expenditure in the semi-monthly (Column 1) and adjusted monthly (Column 2) panel datasets. Both columns show that the null hypothesis of unit roots in all the panels is rejected at the conventional level. The results of the unit root tests for consumption subcategories and variables used to test the risk-coping mechanisms are not reported in this table because those are materially similar results. In Section 4, I used the cluster-robust variance-covariance estimator in the regressions to allow arbitrary autocorrelations (Arellano 1987). Despite this, I confirmed that the null hypothesis of no serial correlation is not rejected at the conventional level in Wooldridge’s autocorrelation test for almost of the consumption and net income variables (Wooldridge 2002).
@@
Test statistics (1) Semi-Monthly (2) Adjusted Monthly
$P$-statistic $p$-value 0.0000 0.0001
$Z$-statistic $p$-value 0.0000 0.0000
@@
Notes: The results of the Fisher-type panel unit root tests based on augmented Dickey-Fuller (ADF) tests are reported in this table. The null hypothesis is that all the panels contain unit roots, whereas the alternative hypothesis is that at least one panel is stationary. In all the specifications, the process under the null hypothesis is assumed to be a random walk with drift. The demeaned data are used to address the effect of cross-sectional dependence. The number of lagged differences in the ADF regression equation reported is set as one because including two or more two lags reduces a large part of the panel units available for computing the test statistics due to the lack of balance of the panels. I confirm that including two lags in the semi-monthly panels does not change the results. Including two lags in the adjusted monthly panels and three lags in the semi-monthly panels is impossible. See Choi (2001) for the details of the tests. These tests need both the first-differenced dependent variable and the lagged dependent variable. Therefore, the units with less than four time-series observations are excluded in the regressions.

### v14

@@ B.6.2 Ancillary Wage
Notes: This scatter plot shows the relationship between the log-transformed number of factory workers and the log-transformed share of ancillary wage. The raw data are listed in Table B.3. The dashed line indicates the linear fit using the log-transformed variables: the fitted regression is $\hat{y}=0.747+0.428x$ with $R^{2}=0.82$. The predictions are materially similar if I use the fractional polynomial using the raw data on the number of factory workers and share of ancillary wage in Table B.3. In the case of fitting based on the raw data, the fitted fractional polynomial model is $\hat{y}^{\textit{Daily Wage}}=53.6-399.3a^{-1/2}+(2.84\times 10^{-10})a^{3}$ with $R^{2}=0.87$. I prefer to use linear regression herein because it can provide more precise fitted values for small-and medium-scale factories. Source: Kitazawa (1924, pp. 20; 58–62).
The daily wage estimated in Section B.6.1 is a fixed salary and thus, does not include additional salaries such as bonuses and allowances (hereafter, wages other than regular salaries are collectively referred to as ancillary wages). However, there are no systematic statistics on the ancillary wages throughout the prewar period (Committee for the History of the Labor Movement 1964, pp. 14–15). Moreover, since the Tsukishima Survey targets skilled worker households, information on the ancillary wages paid to skilled workers is necessary. Fortunately, I found a survey report named Wage Survey Report edited by Kitazawa (1924). This report includes the result of a survey on the wages for machine finishers (fitters) and turners at six machine factories in Tokyo from May–October 1921. Thus, this is particularly useful material to know the wage structure of skilled factory workers.
@@
Panel A in Table B.3 summarizes the survey results. The surveyed factories included $86$, $157$, $376$, $623$, $928$, and $2,837$ workers, ranging from small to large plants. The percentage of skilled workers (finishers and turners) in the surveyed factories ranged from 2727% to 7777%, and the portion of monthly ancillary wages was 12–53%. Figure B.9 is a linear regression fit using log-transformed data on the percentage of ancillary wages and the number of workers. There is a positive correlation between the factory size and ancillary wages. This is consistent with the historical fact that larger factories have more extensive benefits than smaller ones (Odaka 1999).
@@
Panel B of Table B.3 summarizes the distributions of the number of factories, the total number of workers, and the estimated average share of ancillary wage in Tokyo City (Column 1) and Tsukishima (Column 2). The average ancillary wage at the Tsukishima factories is 1212%, which is slightly larger than the value for the entire city of Tokyo (99%). The percentage shares of the number of factory workers by each bin are listed in Columns (1) and (2) of Panel C in Table B.3. As shown, Tsukishima factories tended to have larger factories than those of the entire city, which is reflected in the slightly larger estimate in the average ancillary wage for the Tsukishima factories. In Column (3) of Panel C in Table B.3, the number of THBS heads whose occupation, factory name, and factory size are available are listed. While the percentage shares of the heads in each bin show a similar distribution to that of the Tsukishima factories, the share of middle (large) sized factories is larger (smaller) than the population. If interpreted literally, this tendency implies that the (unobservable) true average ancillary wage for the THBS heads may be slightly lower than the population mean. However, note that the purpose of Panel C in Table 2 is to compare the average monthly wage of the skilled factory workers employed in the Tsukishima factories to the THBS heads’ average monthly earnings. Therefore, the purpose herein is to estimate the average ancillary wage for the target population, for example, the Tsukishima factories, but not to calculate the average ancillary wage for the THBS sample itself.
@@
Notes: Panel A: This panel shows the size of factories and the share of fixed and ancillary monthly wages in six factories surveyed in Kitazawa (1924). The number of factory workers surveyed in factories A–F are 14, 23, 9, 11, 11, and 16, respectively. The share of these surveyed skilled workers to all skilled workers is 20.3 in Factory A, 50.0 in Factory B, 5.3 in Factory C, 2.3 in Factory D, and 4.2 in Factory F, respectively. This figure for factory E is unavailable because the number of skilled workers is unknown (Kitazawa 1924, p. 21). Panel B: Subcolumns 1–2 and 4–5 summarize the number of factories and total number of workers in Tokyo City and Tsukishima, respectively (Handbook of Factories 1921). Subcolumns 3 and 6 show the average share of ancillary monthly wage predicted in Figure B.9 for Tokyo City and Tsukishima, respectively. Panel C: Columns 1–3 summarize the number of factory workers and its share in percentage points in Tokyo City, Tsukishima, THBS sample, respectively. The statistics in Columns 1 and 2 (same figures used in Panel B) are from the Handbook of Factories (1921). In Column 3, the THBS heads who worked in the machinery factories are considered. There are 28 heads whose factory names and size are available. For a blacksmith included in these heads, I assume the number of workers to be ten, based on the report (Department of Health, Ministry of the Interior 1923a, p. 412). Sources: Panel A: Kitazawa (1924, pp. 20; 58–62). Panel B: The number of factory workers is taken from the Ministry of Agriculture and Commerce (1921, pp. 613–629; 674). The average shares of ancillary wage are calculated using Kitazawa (1924, pp. 20; 58–62). See also Figure B.9. Panel C: Data on the number of factory workers and factories are from the Ministry of Agriculture and Commerce (1921, pp. 613–629; 674). Data on the number of heads are from the THBS sample.
@@ B.6.3 Estimating Monthly Income of Adult Male Factory Workers
Panel B: Estimated Monthly Earning: $w^{Total}$
@@
Notes: 1. Average daily wage ($w^{Daily}_{j}$): The average daily wage of male factory workers in the skilled age range (aged 30–40) in the machinery sector is calculated using two official reports of the manufacturing censuses documented in the ASCT and LST. Details in the calculation steps are summarized in Online Appendix B.6.1.
@@
where $w^{\textit{Regular}}>0$ is the monthly regular wage and $w^{\textit{Ancillary}}\geq 0$ indicates the monthly ancillary wage. Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$. [^84] Strictly, this weight may not be required to be in this range. I use this range because all the cases reported in Kitazawa (1924) take the weights between zero and one. The monthly wage equation is then specified as follows:
@@ B.7 Measurement Errors in the Unadjusted Panel Dataset
Note: Figures B.10aB.10cB.10e show the average semi-monthly income and expenditure calculated from unadjusted, shifted, and adjusted panel datasets, respectively. Unadjusted series defines the days from the 1st to the 15th of each month as the first half, whereas the days from the 16th to the end of each month as the second half. The shifted series defines the days from the 2nd to the 16th of each month as the first half, whereas the days from the 17th to the 1st of the following month are the second half. Adjusted series is the baseline definition in the main text, which uses the days from 12th to 26th in each month as the “first-half” and those from 27th to 11th in the next month as the “second-half.” As explained in Section 3, the raw series contains considerable measurement errors due to the timing of payday. The shifted series can manage the measurement errors from the timing of payday but cannot adjust the difference in the payday timing and expenses. In other words, the shifted series ignores the households’ decision on the consumption schedule in each semi-month. The adjusted series reflecting the exact timing of the household expenses can manage measurement errors. Figures B.10bB.10dB.10f illustrate the average semi-monthly income minus expenditure calculated from unadjusted, shifted, and adjusted panel datasets, respectively. Source: Created by the author using the THBS sample.
@@ B.8 Testing Attenuation Bias in the Unadjusted Panel Dataset
I assess the validity of the aggregation method used in this study. Adjusting the aggregate data to account for the timing of paydays reveals more transparent relationships between income and expenditure than the unadjusted data (Section 3.3). To demonstrate this mechanism, I listed the estimates from the unadjusted monthly panel dataset in Column (2) of Table B.5. The estimate for total consumption is $0.26$, which is much lower than the estimate for the adjusted series ($0.43$) in Column (1). In addition, most subcategories show systematically smaller estimates than those from the adjusted data listed in Column (2). For example, the estimate for the food category is now $0.10$ and statistically insignificant. This means that measurement errors in the unadjusted aggregation lead to strong attenuation bias, which results in false negatives (i.e., false rejections of the null of full risk-sharing hypothesis) in most of the consumption categories. I, therefore, specified a generating process of systematic measurement errors in the aggregate survey data, which has attracted attention in the macroeconomics literature (Cochrane 1991; Dynarski et al. 1997; Gervais and Klein 2010; Nelson 1994). Moreover, several subcategories select values close to one or large negative values, which are no longer economically interpretable. [^85] For example, the estimate for housing was negative. The households paid a fixed rent amount monthly. Given that the rent payday was fixed, the shifts in wage paydays might create cells with less income with a fixed amount of rent. In the case of seasonal consumption categories, the shifts may coincide with the timing of seasonal spending. The THBS households tended to purchase work clothes in March to prepare for the start of the new fiscal year in April. They also spent more on clothing in June, likely to buy summer clothes. The shift of paydays to these months may have created a strong but false positive correlation. This means that the miss-assignment of paydays in the unadjusted data may cause critical statistical inference issues in a few subcategories.
@@ B.9 Considering Child Share as a Preference Shock Variable
As explained in the main text, the child share variable exhibits more within-variations than that of the family size. Table B.6 presents the results from specifications that include the share of children under the primary school age range, specifically those aged 0–5, as an alternative preference shock variable. Panels A and B show the results for the consumption smoothing and risk-coping strategy regressions, respectively. The results are materially similar to the main results (Tables 4 and 5). Note that the results remain robust to different definitions of age bins (not reported).
@@
(1) Semi-monthly Panels (2) Monthly Panels
@@
(1) Semi-monthly Panels (2) Monthly Panels
@@
Other members’ earnings 0.001 [0.010] 289 0.036 [0.033] 124
@@
Notes: Panel A: This table presents the results of equation 1. It reports regressions of 11 measures of log-transformed consumption on log-transformed disposable income, the family size control, the household fixed effect, and the time fixed effect. Columns 1 and 2 show the results for the adjusted semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on log-transformed disposable income are shown in the columns labeled “Coef.” Panel B: This table shows the results of equation 2. The proportion of children aged 0–5, household fixed effects, and time-fixed effects are included in all the regressions. Columns 1 and 2 show the estimates for the adjusted semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on the head’s earnings are listed in the “Coef.” columns. Panel B-1: Each net income variable is defined as income minus expenditure. “Net savings”: withdrawal amount minus deposits to savings. “Net insurance”: insurance payments received minus the expenses paid for insurance. “Net borrowing”: the amount borrowed minus debt payments. This category includes net income from both pawnshop and other lending institutions. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts”: total gifts received, including both pecuniary and non-pecuniary, minus any payments for gifts. Panel B-2: “Other family members’ earnings”: total income earned by all family members except the household head. “Sales of miscellaneous assets”: include the sale of goods such as old newspapers and empty bottles. All dependent variables are in yen.
@@ B.10 Distribution of the Idiosyncratic Shocks
Notes: This figure shows the distribution of idiosyncratic shocks in the adjusted semi-monthly THBS panel dataset. The idiosyncratic shocks are estimated as the residuals from the regression of the first difference in income on the first difference in aggregate income. The red line indicates the mean of the idiosyncratic shocks (0.235-0.235). Source: Created by the author using the THBS sample.
Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset. [^86] I use the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets. This is because the daily panels include a set of consecutive household-date cells with no income. These cells cannot be used to predict the idiosyncratic shocks as the residuals are economically meaningfully, and thus it is practically impossible to illustrate the idiosyncratic shocks based on the daily dataset. The idiosyncratic shocks are calculated as residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919. For household $i$ in semi-month $t$ cell, the residual is then defined as follows:
@@ B.11 Unit Root and Autocorrelations
Table B.7 presents the results of the Fisher-type panel unit root tests for the total expenditure in the semi-monthly (Column 1) and monthly (Column 2) panel datasets. Both columns show that the null hypothesis of unit roots in all the panels is rejected at the conventional level. The results of the unit root tests for consumption subcategories and variables used to test the risk-coping mechanisms are not reported in this table because those are materially similar results. In Section 4, I used the cluster-robust variance-covariance estimator in the regressions to allow arbitrary autocorrelations (Arellano 1987). Despite this, I confirmed that the null hypothesis of no serial correlation is not rejected at the conventional level in Wooldridge’s autocorrelation test for almost of the consumption and net income variables (Wooldridge 2002).
@@
Test statistics (1) Semi-Monthly Panels (2) Monthly Panels
$P$-statistic pp-value 0.0000 0.0001
$Z$-statistic pp-value 0.0000 0.0000
@@
Notes: The results of the Fisher-type panel unit root tests based on augmented Dickey-Fuller (ADF) tests are reported in this table. The null hypothesis is that all the panels contain unit roots, whereas the alternative hypothesis is that at least one panel is stationary. In all the specifications, the process under the null hypothesis is assumed to be a random walk with drift. The demeaned data are used to address the effect of cross-sectional dependence. The number of lagged differences in the ADF regression equation reported is set as one because including two or more two lags reduces a large part of the panel units available for computing the test statistics due to the lack of balance of the panels. I confirm that including two lags in the adjusted semi-monthly panels does not change the results. Including two lags in the adjusted monthly panels and three lags in the adjusted semi-monthly panels is impossible. See Choi (2001) for the details of the tests. These tests need both the first-differenced dependent variable and the lagged dependent variable. Therefore, the units with less than four time-series observations are excluded in the regressions.